# MambaTAD: When State-Space Models Meet Long-Range Temporal Action Detection

### v1

@@ I Introduction
Temporal action detection (TAD) aims to detect specific action categories and extract corresponding temporal spans in untrimmed videos. It is a long-standing and challenging problem in video understanding with extensive real-world applications such as sports analysis, surveillance and security. The development of deep neural networks such as CNNs [shi2023tridet, yang2024dyfadet] and Transformers [tan2021relaxed, yang2024adapting] has led to continuous advancements in TAD performance over the past few years. However, CNNs have limited capabilities in capturing long-range dependencies, while Transformers face challenges with computational complexity and feature discrimination [shi2023tridet].
@@
Recently, Structured State-Space Sequence models (S4) [gu2021efficiently] such as Mamba [gu2023mamba] have demonstrated great efficiency and effectiveness in deep network construction [DBLP:conf/icml/ZhuL0W0W24, ma2024u]. These models, enhanced by specially designed structured re-parameterization [gu2020hippo] and selective scan mechanisms, facilitate the natural activation of extended temporal moments, thereby improving classification and boundary regression performance. However, the standard Mamba, which is designed for long sequence data in natural language using one forward branch, is not a natural fit for the TAD task. Specifically, Mamba processes flattened one-dimensional sequences in a recursive manner. It often loses temporal information of earlier moments and suffers from the problem of decay of temporal information due to the involved lower triangular matrices [dao2024transformers]. In addition, since the trainable weights are the incorporation of a lower triangular matrix and an upper triangular matrix in bidirectional Mamba [li2024videomamba, chen2024video], they often face the problem of self-element conflict (i.e., diagonal conflict) while modeling global visual context. Such self-element conflict impedes the model from learning critical temporal boundary representations effectively.
@@
These observations motivate us to present MambaTAD, a novel state-space model for effectively handling the challenging TAD problem. MambaTAD consists of a detector including two novel designs and an adapter for the backbone. The first is a pyramid model featuring a Diagonal-Masked Bidirectional State-Space (DMBSS) module. The module introduces a dual-branch input inversion mechanism that mitigates the decay of temporal information better and preserves global temporal information by retrieving earlier information via input flipping. This directly helps capture the progression and structure of long-duration actions. On top of that, the model mitigates diagonal conflict by masking the diagonal elements in backward matrices. The second is a global feature fusion head that integrates features of different granularities to enhance global awareness, enabling the model to capture both fine-grained details and broad patterns like slow motion in actions. At the instance level, it means that the model can recognize small, fast movements as well as the overall structure and flow of the action across time. As in Figure 1, MambaTAD outperforms DyFADet [yang2024dyfadet] by 11.2% and 7.0% in mAP for long-duration actions under the Coverage and Length categories, respectively.
@@
What is more, end-to-end training in TAD involves simultaneously optimizing both the video encoder and the action detector as a unified model. To further extend the capability of DMBSS beyond the detector, we propose the State-Space Temporal Adapter (SSTA), a novel parameter-efficient module designed for end-to-end TAD. Unlike conventional fine-tuning approaches that require extensive parameter updates[zhao2023re2tal, yang2024adapting], SSTA leverages DMBSS within a state-space modeling framework, enabling efficient adaptation while preserving global and local temporal dependencies. Extensive experiments over five TAD datasets show that MambaTAD outperforms SOTA methods consistently with fewer parameters and lower FLOPs.
@@ II-A Temporal Action Detection
TAD methods are broadly classified as two-stage and one-stage approaches. Two-stage methods [qing2021temporal, lin2019bmn, liu2019multi] involve proposal generation and classification but face challenges due to high complexity, limiting end-to-end training. Recent trends favor one-stage frameworks, which support end-to-end optimization. One-stage methods are further divided into anchor-based and anchor-free approaches. Anchor-based methods [chen2019relation, chen2020afnet, lin2017single, long2019gaussian] assign multi-scale intervals to temporal locations, while anchor-free methods [lin2021learning, xia2023exploring, zhao2022temporal] generate flexible proposals with precise boundaries. MambaTAD falls into the one-stage anchor-free category. Different methods employ CNNs [yang2023basictad, yang2020revisiting, liu2018deep, yan2020higcin], Graph models [zhao2021video, zeng2019graph], Transformers [liu2022end, 10440542, yan2023progressive, su2023sequence], or Diffusion models [nag2023difftad] for temporal encoding, whereas our approach is based on state-space models.
TAD methods are also classified into feature-based and end-to-end approaches. Feature-based methods use pre-extracted RGB and optionally optical flow features, while end-to-end methods process raw video frames, jointly optimizing the video encoder and action detector[liu2020progressive]. E2E-TAD [liu2022empirical] analyzed design choices affecting TAD performance and proposed a mid-resolution baseline due to GPU constraints. Advancements like TALLFormer [cheng2022tallformer], which introduced partial data backpropagation, and Re2TAL [zhao2023re2tal], which leveraged reversible architectures for memory efficiency, have much improved TAD. AdaTAD [liu2024end] further scaled TAD to 1 billion parameters via PEFT. To our knowledge, we are the first to explore end-to-end TAD in state-space models.
@@ II-B State-Space Models for Long-range Modeling
State-Space Models (SSMs) have gained attention for their efficient linear scaling in long-range dependency modeling. The Structured State-Space Sequence model [gu2021efficiently] pioneered this area, followed by the S5 layer [smith2022simplified], which introduced MIMO SSM and parallel scanning. Advances like H3 [fu2022hungry] narrowed the gap with Transformers, while the Gated State Space layer [mehta2022long] enhanced S4 with gating. Recently, Mamba [gu2023mamba] introduced a selective mechanism and optimized hardware design, surpassing Transformers in large-scale tasks while preserving linear complexity. Mamba2 [dao2024transformers] further improves via viewing Transformers as structured state space models, enabling efficient algorithms via state space duality, while Hydra [hwang2024hydra] uses bidirectional SSMs with matrix mixers to better capture past and future context. Our approach is developed to tackle the challenges of applying Mamba in TAD and, in principle, can be extended to a wide range of foundational state-space models.
@@ II-C Visual State-Space Models
Several studies have applied SSMs to various visual foundation models or developed hybrid architectures by combining SSMs with convolution or attention mechanisms. For instance, [DBLP:conf/icml/ZhuL0W0W24, li2024videomamba, lu2024videomambapro] integrate bidirectional SSMs for data-dependent global visual context modeling and employ position embedding to enhance location-aware visual understanding. Moreover, many pioneering works have applied visual state-space models to visual tasks, like [islam2022long] utilized 1D S4 to manage long-range temporal dependencies for video classification, [ma2024u] proposes a hybrid CNN-SSM architecture to handle the long-range dependencies for biomedical image segmentation, [mondal2024hummuss] for human motion understanding, and processes human motion using SSMs, [guo2024mambair] explores the potential of Mamba to for image restoration. Recently, CausalTAD [liu2024harnessing] further leverages temporal causality in SSMs to improve temporal action detection by integrating with causal attention blocks, while our DMBSS remains purely SSM-based and lightweight. [chen2024video] proposed a Decomposed Bidirectionally Mamba (DBM) block for many video understanding tasks, including temporal action detection. Our approach further addresses the limitations of Mamba in TAD while introducing the DMBSS and a global awareness detection head, enhancing flexibility and efficiency in modeling both action and boundary representations.
@@ III-A Preliminaries
where the variable $\bar{\mathbf{K}}\in\mathbb{R}^{s}$ corresponds to a structured convolution kernel derived from SSMs, $s$ is the length of the input sequence, and $\ast$ is the convolution operation. The cutting-edge state-space model, Mamba [gu2023mamba] proposed a selective scan mechanism, which refined elements $\mathbf{B,C}$ and $\Delta$ to become input-adaptive, thereby enabling a more dynamic feature representation:
@@
Problem definition. MambaTAD is a unified end-to-end one-stage anchor-free and attention-free architecture that processes raw untrimmed videos or temporal video features as input and produces representations for each action category and their temporal localization. Let the input of MambaTAD be videos $\mathbf{X}$, consisting of a series of frames or video clips features extracted at different time steps $t$ by pre-trained models [carreira2017quo, alwassel2021tsp, tong2022videomae, wang2023videomae, wang2024internvideo2], represented as $\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{t},\ldots,\mathbf{x}_{T}\}$, where the total duration $T$ is different across videos. For simplicity, $x_{t}$ and $\mathbf{x}_{t}$ both denote the $t$-th discrete token, where $\mathbf{x}_{t}\in\mathbb{R}^{c}$ in video feature representation. MambaTAD learns to model the underlying discrete signal resulting from videos and the action label $\mathbf{Y}=\{\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{I}\}$, where $I$ is the number of ground truth labels in a video and $\mathbf{y}_{i}=(start_{i},end_{i},action_{i})$ to produce action proposals $(start_{i},end_{i})$ and categories $action_{i}$ during inference. Here, $start_{i},end_{i}\in[0,T]$ and $start_{i}<end_{i}$, $action_{i}\in\{0,C-1\}$ where $C$ is pre-defined categories.
@@ III-B The Framework
The framework applies a Parameter-efficient fine-tuning (PEFT) mechanism to fine-tune a plug-and-play module named State-Space Temporal Adapter (SSTA) to achieve efficient and effective transfer learning for end-to-end temporal action detection and build a simple yet effective anchor-free action detector. As shown in Fig. 2, MambaTAD consists of three main parts: a video backbone, a projection pyramid with Diagonal-masked Bidirectional State-Space (DMBSS) modules, and a global feature fusion head. We begin by leveraging video backbones [carreira2017quo, alwassel2021tsp, tong2022videomae, wang2023videomae, wang2024internvideo2] to extract spatiotemporal features. To efficiently model long-range dependencies in an end-to-end manner, we introduce SSTA that captures both local and global non-causal information with a smaller parameter overhead. It is worth noting that we consider both end-to-end (E2E) and non-end-to-end settings, while SSTA is utilized exclusively in the end-to-end setting. Following that, we construct pyramid projection layers incorporating the proposed DMBSS to model action instances across varying temporal durations. While recent TAD methods [zhang2022actionformer, shi2023tridet, gan2022temporal] iteratively downsample temporal features, our approach uniquely applies DMBSS at each scale to enhance cross-scale feature interactions. Lastly, we fuse multi-granularity representations from different pyramid levels to improve global contextual understanding, leading to more precise action detection. We will elaborate on the proposed modules as follows.
@@ III-C Diagonal-Masked Bidirectional State-Space Module
Given an input $\mathbf{X}\in\mathbb{R}^{s\times c}$ where $s$ denotes the sequence length after feature sampling, proportional to the original video duration $T$, and $c$ is the feature channel, the output $\mathbf{X}_{out}$ maintains the same dimensional structure as the input. Mamba [gu2023mamba] is originally designed for the 1D sequence under causal relations in language tasks. Following [dao2024transformers], the final output in Eq. 4 can be simplified as:
@@
From $\mathbf{M}$, it is evident that the earlier state tends to exhibit more zero weights, a phenomenon referred to as causality. In NLP and text generation, causality is fundamental [hu2021causal], as autoregressive models generate text sequentially, relying solely on preceding tokens to maintain logical coherence and prevent future information from influencing past decisions. However, in the TAD task, a video sequence’s future state depends not only on prior tokens but also on the broader temporal context, where identifying action boundaries requires information from both preceding and succeeding tokens. As a result, enforcing strict causality leads to temporal information loss by restricting the model’s ability to incorporate future context for more precise action localization. On the other hand, bidirectional temporal information plays a key role in video representation modeling. [zhu2017bidirectional] exploits bidirectional processing in combination with multirate feature reconstruction to integrate both past and future context across multiple temporal resolutions, effectively mitigating information loss over long sequences. Similarly, a Decomposed Bidirectionally Mamba block (DBM)[chen2024video] applies a Mamba [gu2023mamba] module with shared parameters for bidirectional scanning, allowing more effective temporal feature processing, which partially mitigates this issue. Specifically, let the feature sequence of length $s$ with channel dimension $c$ be concatenated as:
@@
As illustrated in Alg. 1, we expand the feature dimensionality to $\lambda c$ (with $\lambda=4$) via a linear transformation after LayerNorm, enabling dual-branch bidirectional processing. The expanded representation is split into two equal parts for forward and backward streams. To enforce bidirectional parameter sharing [chen2024video], the second segment is inverted before concatenation. Besides, we propose a dual-branch approach, without parameter sharing, that enhances temporal modeling and mitigates the decay of temporal information via two separate branches. Our ablation study in Sec. IV-F gives experimental analysis for this. To mitigate the issue of diagonal conflict, we introduce a masking mechanism that nullifies the diagonal elements in the trainable weights for each backward branch. The $SSM(\cdot)$ in Alg. 1 stands for vanilla mamba [gu2023mamba] without output linear layer, thereby reducing the output dimensionality from $2c$ to $c$. That is because the application of a final linear transformation at the end of the dual-branch processing enhances dual-branch feature fusion. After diagonal masking, bidirectional dual branches integrate to capture richer temporal contexts. The representations are split along the batch dimension and then merged along the channel dimension for consolidation. Lastly, a final linear layer integrates both temporal and channel information from two separate branches and restores the original channel size, yielding the output token.
@@
where $flip(\cdot)$ is flipping operation along the sequence dimension, $DWConv(\cdot)$ represents depth-wise convolution, $SSM(\cdot)$ is selective scan mechanism  [gu2023mamba] without out linear layer and $\sigma(\cdot)$ denotes SiLU activation [shazeer2020glu]. Following  [chen2024video], features from the three pathways are aggregated with element-wise product and concatenation:

### v2

@@ I Introduction
Temporal action detection (TAD) aims to detect specific action categories and extract corresponding temporal spans in untrimmed videos. It is a long-standing and challenging problem in video understanding with extensive real-world applications such as sports analysis, surveillance and security. The development of deep neural networks such as CNNs [62, 80] and Transformers [67, 82] has led to continuous advancements in TAD performance over the past few years. However, CNNs have limited capabilities in capturing long-range dependencies, while Transformers face challenges with computational complexity and feature discrimination [62].
@@
Recently, Structured State-Space Sequence models (S4) [16] such as Mamba [15] have demonstrated great efficiency and effectiveness in deep network construction [97, 54]. These models, enhanced by specially designed structured re-parameterization [14] and selective scan mechanisms, facilitate the natural activation of extended temporal moments, thereby improving classification and boundary regression performance. However, the standard Mamba, which is designed for long sequence data in natural language using one forward branch, is not a natural fit for the TAD task. Specifically, Mamba processes flattened one-dimensional sequences in a recursive manner. It often loses temporal information of earlier moments and suffers from the problem of decay of temporal information due to the involved lower triangular matrices [10]. In addition, since the trainable weights are the incorporation of a lower triangular matrix and an upper triangular matrix in bidirectional Mamba [27, 6], they often face the problem of self-element conflict (i.e., diagonal conflict) while modeling global visual context. Such self-element conflict impedes the model from learning critical temporal boundary representations effectively.
@@
These observations motivate us to present MambaTAD, a novel state-space model for effectively handling the challenging TAD problem. MambaTAD consists of a detector including two novel designs and an adapter for the backbone. The first is a pyramid model featuring a Diagonal-Masked Bidirectional State-Space (DMBSS) module. The module introduces a dual-branch input inversion mechanism that mitigates the decay of temporal information better and preserves global temporal information by retrieving earlier information via input flipping. This directly helps capture the progression and structure of long-duration actions. On top of that, the model mitigates diagonal conflict by masking the diagonal elements in backward matrices. The second is a global feature fusion head that integrates features of different granularities to enhance global awareness, enabling the model to capture both fine-grained details and broad patterns like slow motion in actions. At the instance level, it means that the model can recognize small, fast movements as well as the overall structure and flow of the action across time. As in Figure 1, MambaTAD outperforms DyFADet [80] by 11.2% and 7.0% in mAP for long-duration actions under the Coverage and Length categories, respectively.
@@
What is more, end-to-end training in TAD involves simultaneously optimizing both the video encoder and the action detector as a unified model. To further extend the capability of DMBSS beyond the detector, we propose the State-Space Temporal Adapter (SSTA), a novel parameter-efficient module designed for end-to-end TAD. Unlike conventional fine-tuning approaches that require extensive parameter updates[93, 82], SSTA leverages DMBSS within a state-space modeling framework, enabling efficient adaptation while preserving global and local temporal dependencies. Extensive experiments over five TAD datasets show that MambaTAD outperforms SOTA methods consistently with fewer parameters and lower FLOPs.
@@ II-A Temporal Action Detection
TAD methods are broadly classified as two-stage and one-stage approaches. Two-stage methods [59, 36, 40] involve proposal generation and classification but face challenges due to high complexity, limiting end-to-end training. Recent trends favor one-stage frameworks, which support end-to-end optimization. One-stage methods are further divided into anchor-based and anchor-free approaches. Anchor-based methods [7, 5, 37, 49] assign multi-scale intervals to temporal locations, while anchor-free methods [35, 76, 96] generate flexible proposals with precise boundaries. MambaTAD falls into the one-stage anchor-free category. Different methods employ CNNs [81, 79, 41, 78], Graph models [94, 90], Transformers [47, 31, 77, 66], or Diffusion models [58] for temporal encoding, whereas our approach is based on state-space models.
TAD methods are also classified into feature-based and end-to-end approaches. Feature-based methods use pre-extracted RGB and optionally optical flow features, while end-to-end methods process raw video frames, jointly optimizing the video encoder and action detector[42]. E2E-TAD [46] analyzed design choices affecting TAD performance and proposed a mid-resolution baseline due to GPU constraints. Advancements like TALLFormer [8], which introduced partial data backpropagation, and Re2TAL [93], which leveraged reversible architectures for memory efficiency, have much improved TAD. AdaTAD [45] further scaled TAD to 1 billion parameters via PEFT. To our knowledge, we are the first to explore end-to-end TAD in state-space models.
@@ II-B State-Space Models for Long-range Modeling
State-Space Models (SSMs) have gained attention for their efficient linear scaling in long-range dependency modeling. The Structured State-Space Sequence model [16] pioneered this area, followed by the S5 layer [63], which introduced MIMO SSM and parallel scanning. Advances like H3 [12] narrowed the gap with Transformers, while the Gated State Space layer [56] enhanced S4 with gating. Recently, Mamba [15] introduced a selective mechanism and optimized hardware design, surpassing Transformers in large-scale tasks while preserving linear complexity. Mamba2 [10] further improves via viewing Transformers as structured state space models, enabling efficient algorithms via state space duality, while Hydra [24] uses bidirectional SSMs with matrix mixers to better capture past and future context. Our approach is developed to tackle the challenges of applying Mamba in TAD and, in principle, can be extended to a wide range of foundational state-space models.
@@ II-C Visual State-Space Models
Several studies have applied SSMs to various visual foundation models or developed hybrid architectures by combining SSMs with convolution or attention mechanisms. For instance, [97, 27, 51] integrate bidirectional SSMs for data-dependent global visual context modeling and employ position embedding to enhance location-aware visual understanding. Moreover, many pioneering works have applied visual state-space models to visual tasks, like [25] utilized 1D S4 to manage long-range temporal dependencies for video classification, [54] proposes a hybrid CNN-SSM architecture to handle the long-range dependencies for biomedical image segmentation, [57] for human motion understanding, and processes human motion using SSMs, [17] explores the potential of Mamba to for image restoration. Recently, CausalTAD [44] further leverages temporal causality in SSMs to improve temporal action detection by integrating with causal attention blocks, while our DMBSS remains purely SSM-based and lightweight. [6] proposed a Decomposed Bidirectionally Mamba (DBM) block for many video understanding tasks, including temporal action detection. Our approach further addresses the limitations of Mamba in TAD while introducing the DMBSS and a global awareness detection head, enhancing flexibility and efficiency in modeling both action and boundary representations.
@@ III-A Preliminaries
where the variable $\bar{\mathbf{K}}\in\mathbb{R}^{s}$ corresponds to a structured convolution kernel derived from SSMs, $s$ is the length of the input sequence, and $\ast$ is the convolution operation. The cutting-edge state-space model, Mamba [15] proposed a selective scan mechanism, which refined elements $\mathbf{B,C}$ and $\Delta$ to become input-adaptive, thereby enabling a more dynamic feature representation:
@@
Problem definition. MambaTAD is a unified end-to-end one-stage anchor-free and attention-free architecture that processes raw untrimmed videos or temporal video features as input and produces representations for each action category and their temporal localization. Let the input of MambaTAD be videos $\mathbf{X}$, consisting of a series of frames or video clips features extracted at different time steps $t$ by pre-trained models [4, 1, 71, 73, 74], represented as $\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{t},\ldots,\mathbf{x}_{T}\}$, where the total duration $T$ is different across videos. For simplicity, $x_{t}$ and $\mathbf{x}_{t}$ both denote the $t$-th discrete token, where $\mathbf{x}_{t}\in\mathbb{R}^{c}$ in video feature representation. MambaTAD learns to model the underlying discrete signal resulting from videos and the action label $\mathbf{Y}=\{\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{I}\}$, where $I$ is the number of ground truth labels in a video and $\mathbf{y}_{i}=(start_{i},end_{i},action_{i})$ to produce action proposals $(start_{i},end_{i})$ and categories $action_{i}$ during inference. Here, $start_{i},end_{i}\in[0,T]$ and $start_{i}<end_{i}$, $action_{i}\in\{0,C-1\}$ where $C$ is pre-defined categories.
@@ III-B The Framework
The framework applies a Parameter-efficient fine-tuning (PEFT) mechanism to fine-tune a plug-and-play module named State-Space Temporal Adapter (SSTA) to achieve efficient and effective transfer learning for end-to-end temporal action detection and build a simple yet effective anchor-free action detector. As shown in Fig. 2, MambaTAD consists of three main parts: a video backbone, a projection pyramid with Diagonal-masked Bidirectional State-Space (DMBSS) modules, and a global feature fusion head. We begin by leveraging video backbones [4, 1, 71, 73, 74] to extract spatiotemporal features. To efficiently model long-range dependencies in an end-to-end manner, we introduce SSTA that captures both local and global non-causal information with a smaller parameter overhead. It is worth noting that we consider both end-to-end (E2E) and non-end-to-end settings, while SSTA is utilized exclusively in the end-to-end setting. Following that, we construct pyramid projection layers incorporating the proposed DMBSS to model action instances across varying temporal durations. While recent TAD methods [91, 62, 13] iteratively downsample temporal features, our approach uniquely applies DMBSS at each scale to enhance cross-scale feature interactions. Lastly, we fuse multi-granularity representations from different pyramid levels to improve global contextual understanding, leading to more precise action detection. We will elaborate on the proposed modules as follows.
@@ III-C Diagonal-Masked Bidirectional State-Space Module
Given an input $\mathbf{X}\in\mathbb{R}^{s\times c}$ where $s$ denotes the sequence length after feature sampling, proportional to the original video duration $T$, and $c$ is the feature channel, the output $\mathbf{X}_{out}$ maintains the same dimensional structure as the input. Mamba [15] is originally designed for the 1D sequence under causal relations in language tasks. Following [10], the final output in Eq. 4 can be simplified as:
@@
From $\mathbf{M}$, it is evident that the earlier state tends to exhibit more zero weights, a phenomenon referred to as causality. In NLP and text generation, causality is fundamental [23], as autoregressive models generate text sequentially, relying solely on preceding tokens to maintain logical coherence and prevent future information from influencing past decisions. However, in the TAD task, a video sequence’s future state depends not only on prior tokens but also on the broader temporal context, where identifying action boundaries requires information from both preceding and succeeding tokens. As a result, enforcing strict causality leads to temporal information loss by restricting the model’s ability to incorporate future context for more precise action localization. On the other hand, bidirectional temporal information plays a key role in video representation modeling. [98] exploits bidirectional processing in combination with multirate feature reconstruction to integrate both past and future context across multiple temporal resolutions, effectively mitigating information loss over long sequences. Similarly, a Decomposed Bidirectionally Mamba block (DBM)[6] applies a Mamba [15] module with shared parameters for bidirectional scanning, allowing more effective temporal feature processing, which partially mitigates this issue. Specifically, let the feature sequence of length $s$ with channel dimension $c$ be concatenated as:
@@
As illustrated in Alg. 1, we expand the feature dimensionality to $\lambda c$ (with $\lambda=4$) via a linear transformation after LayerNorm, enabling dual-branch bidirectional processing. The expanded representation is split into two equal parts for forward and backward streams. To enforce bidirectional parameter sharing [6], the second segment is inverted before concatenation. Besides, we propose a dual-branch approach, without parameter sharing, that enhances temporal modeling and mitigates the decay of temporal information via two separate branches. Our ablation study in Sec. IV-F gives experimental analysis for this. To mitigate the issue of diagonal conflict, we introduce a masking mechanism that nullifies the diagonal elements in the trainable weights for each backward branch. The $SSM(\cdot)$ in Alg. 1 stands for vanilla mamba [15] without output linear layer, thereby reducing the output dimensionality from $2c$ to $c$. That is because the application of a final linear transformation at the end of the dual-branch processing enhances dual-branch feature fusion. After diagonal masking, bidirectional dual branches integrate to capture richer temporal contexts. The representations are split along the batch dimension and then merged along the channel dimension for consolidation. Lastly, a final linear layer integrates both temporal and channel information from two separate branches and restores the original channel size, yielding the output token.
@@
where $flip(\cdot)$ is flipping operation along the sequence dimension, $DWConv(\cdot)$ represents depth-wise convolution, $SSM(\cdot)$ is selective scan mechanism  [15] without out linear layer and $\sigma(\cdot)$ denotes SiLU activation [61]. Following  [6], features from the three pathways are aggregated with element-wise product and concatenation: